Calculate the specific charge on alpha particle.. Please explain with details

Grade:11

1 Answers

Tanujit pal
6 years ago
Specific charge= total charge/massfor alpha particle charge is 2e=2*1.6*10^-19 and mass is (2*mass of neutron+2*mass of proton)=(2*1.674*10^-27+2*1.672*^10-27)now e/m for alpha particle is2*1.6*10^-19/2*10^-27(1.674+1.672)=4.78*10^7 C/kg